OpenStudy (anonymous):

simplify using a table (6q + 7)(2q + 3)

OpenStudy (demandandsupply):

Use foil expansion

OpenStudy (demandandsupply):

start with 6q times 2q

OpenStudy (anonymous):

12q

OpenStudy (anonymous):

@DemandAndSupply

OpenStudy (demandandsupply):

6q times 3

OpenStudy (demandandsupply):

and it would be 12q^2

OpenStudy (anonymous):

the answer would be 12q^2 + 32q + 21, right?

OpenStudy (demandandsupply):

Yes

OpenStudy (anonymous):

(6q + 7)(2q + 3) =12q^2 + 18q + 14q + 21 =12q^2 + 32q + 21

OpenStudy (anonymous):

yay!!!, I have one more I need you to check, it's (-4m + 6)(3m + 2)

OpenStudy (demandandsupply):

Sure

OpenStudy (demandandsupply):

Exactly the same procedure as before

OpenStudy (demandandsupply):

start with -4m times 3m

OpenStudy (anonymous):

I think it's -12m^2 + 10m + 12

OpenStudy (demandandsupply):

The middle part is incorrect, you should get -8m +9m

OpenStudy (anonymous):

thats not one of my answer choices

OpenStudy (demandandsupply):

Neverming, i read that wrong, you were right!

OpenStudy (anonymous):

oh, yay!!!!!!!! thx

OpenStudy (demandandsupply):

Apologies

OpenStudy (demandandsupply):

No problem

OpenStudy (anonymous):

it ok

OpenStudy (anonymous):

(-4m + 6)(3m + 2) =-12m^2 - 8m + 18m + 12 =-12m^2 +10m +12 [ANS]